OpenStudy (anonymous):

There Are Many Different Types Of Crystals. This Fact Is Easy To See When They're Magnified , For Example At A Ratio Of 24:1, In CM If The Magnified Crystal Is 12 Cm, How Large Is It In Real Life ?

OpenStudy (anonymous):

give it a try show me step and i will tell you where you get mistakes

OpenStudy (anonymous):

I Dont Know How To Do This One , SORRY :(

OpenStudy (anonymous):

1 cm real life -> 24 cm magnified we can set now :

OpenStudy (anonymous):

\[\frac{ 1 }{ 24 } = \frac{ x }{ 12 }\]

OpenStudy (anonymous):

\[\frac{ 24 }{ 1 }=\frac{ 12 }{ ? }\]

OpenStudy (anonymous):

lol surdawi :)

OpenStudy (anonymous):

use either mine or surdawi's to find x .. you see in my way the "real life" sizes are at the top and magnified are at the bottom in surdawi's is the opposite but both are fine!

OpenStudy (anonymous):

So Im Gonna Cross Multiply

OpenStudy (anonymous):

yse

OpenStudy (anonymous):

yes..

OpenStudy (anonymous):

12*1=12

OpenStudy (anonymous):

where did the x and 24 go

OpenStudy (anonymous):



OpenStudy (anonymous):